quickstudent:

If the area of a rhombus is 126 m^2, and one diagonal measures 18 m, find the length of the other diagonal. 14 cm 7 cm 3.5 cm 72 cm

Allison:

Lol nvm.

Allison:

A. 14 Cm Because the S=(d1*d2)/2 so we have 126=(18*X)/2 or 126*2=18*X X=252/18=14